UT 02 Dec 2021 - DEC time - Split w/AZ Queue



Night Info:

  • Sunset 17:19
  • Twlight Ends 18:40
  • Moon rise 04:49
  • Twlight Begins 05:38
  • Sunrise 07:00

Night Plan:

  • Configure for LUCI-Bino - authorize LUCI-1
  • OSCO - twlight - no NCPA, NCPA on - until 18:40
  • Bin 4 ESM - Eng test J2329-0522 (AZ science target)
    • If conditions and IQ meet PI requirements -AND- AO is stable - stay on for 1hr - counts as AZ time 18:40-19:25
    • If conditions and IQ do NOT meet PI requirements -BUT- AO is stable - take data for ~1hr - counts as ENG time 18:40-19:10
  • Bin 1 ESM - M52 R=10.7 AO Ref - 30-45min
    • /home/lbto/Service_Observing/2021B_D_E/LUCI/AO_ESM/ESM/M52
    • M52_R10pt7_ESM_L1.xml or M52_R15pt2_ESM_SL.xml
  • Bni 2 ESM - M34 R=12.4 (Bin2 target R>12.51 is bin2 crossover - likely this is ok)
    • /home/lbto/Service_Observing/2021B_D_E/LUCI/AO_ESM/ESM/M34
    • M34_R12pt4_ESM_L1.xml or M34_R12pt4_ESM_SL.xml
  • Bin 4 Full AO tests
  • Authorize LUCI-2 only
  • LUCI-2 RotZP and pseudo-Telluric test 19:25-20:10- OR - 19:10-19:55
    • StoneA and Telluric visible until 22:30hr
    • Files: /home/lbto/Service_Observing/Restart_2021/LUCI/RotZP
      • OT File: /home/lbto/Service_Observing/Restart_2021/LUCI/LUCI_2021_Restart_V1.67.xml
      • Files have already been generated for LUCI-2
    • Acquire StoneA mask - check RotZP
      • USE DEFAULT LMS FILE
    • Acquire Telluric star - standard acquisution through StoneA mask
      • use revised LMS file: luci1.138.stnA2020.lms
      • Load the LMS file manually in RTD during acquisition
      • "tricks" OT into thinking several 4"x4" boxes are longslits
    • Take data with Telluric star - script switches to N1.8 camera, but mirror in w/Br Gamma - dithers to several boxes (centered) and slits - verify star lands in slit/box
    • Repeats with the N3.75 camera for finer detail and to rule out any issues with the camera wheel or other settings causing the problem
  • Other DEC tasks - 19:25/19:10 - 21:40

  • If for some reason LUCI-2 RotZP check cannot be done first part of night - M67 check - ~45min total
    • M67 and telluric visible from 00:15-twilight
    • Files: /home/lbto/Service_Observing/Restart_2021/LUCI/RotZP
      • OT File: /home/lbto/Service_Observing/Restart_2021/LUCI/LUCI_2021_Restart_V1.66.xml
      • Files have already been generated for LUCI-2
    • Acquire M67 mask - Check RotZP
      • USE DEFAULT LMS FILE
    • Acquire Telluric star - standard acquisution through M67 mask
      • use the revised LMS file: luci1.139.M67.lms
      • Load the LMS file manually in RTD during acquisition
      • "tricks" OT into thinking several 4"x4" boxes are longslits
    • Take data with Telluric star - script switches to N1.8 camera, but mirror in w/Br Gamma - dithers to several boxes (centered) and slits - verify star lands in slit/box
    • Repeats with the N3.75 camera for finer detail and to rule out any issues with the camera wheel or other settings causing the problem
  • ´╗┐Other DEC tasks -

Summary

0:15 restarted the wfs software. Safed the wfs (ocam cooling off, binnings set, silver mirror in place). w_stop_all/W_start all. House keeper did not come up at first. repeated twice more to get it to come up cleanley. When it came up, set bin1, freq 200, modulation3, gain1 and OCAM live. Set 4x4 binning for the ccd47 and that also live without issue.

0:25 David is opening

0:48 Sending OSCO Preset to HD216399, Intervention mode set.Seeing is 0.65" on the guider. GainsWith NCPA on
luci1.20211202.0021-40

The top image for OSCO looked like the reflected image from ND filter was closer than usual so verified field stop alignment was ok. It looked pretty close.

1:07 Sending preset for ESM bin4 target J2339-0522. Seeing is 0.75" on the guiders. 20211130 TTM table

luci1.202111202.0041 Gains 0.5 to 0.15, 0.827 for slope rms, estimated magnitude on OCAM is 16.9

luci1.20211202.0042 Higher gains. Very elongated image

luci2.20211202.0043 TT1 HO 0.05, still slightly elongated.

obs5 has slowed to a crawl, rebooting it.

1:25 Moving to older TTM table for bin4 and resending preset. The camera lens is railed, but we are going to try and close. Lots of TT jitter. Soon as any HO1 is applied the TT goes bananas. 20211025 table

Gains 1.2, 0,0, luci1..0045

Gains 1.2, 0.05,0 luci1..0046

1:37 Sending another preset with another table, 20211119 table. Pupil centered with lots of range, TT jitter not as bad. When Doug bumps tt the slope RMS actually improves.

Gains 1.15, 0.05 luci1...0047

1:48 Sending a new preset with a new TTM table with pupils that Juan Carlos and Enrico made today. Pupils are not flat. Thin cirrus,

luci1...0048

2:03 Sending preset with frankenstein Pupil. Failed to find pupil file so preset failed.

2:09 Sending again with table corrections made for frankenstein pupils. TT Jitter looks like crap. Bumpin up TT

Gains 1.5 0.05, tt jitter is still pretty high, pupils may be a touch flatter but not much better

luci1.20211202.0049, PSF elongated.

Slope position on displaty looks off. Juan Carlos shifted it so resending preset one last time to see if it resoleved things, thin cirrus, 0.8", thin cirrus.

2:23 Doug is reverting to the TTM table for 20211130 and we are going to move on to TTM bin1 testing.

2:25 Sending Bin 1 ESM - M52 R=10.7 AO Ref, M52_R10pt7_ESM_L1.xml, B8
OCAM estimated magnitude is 11.4

luci1...050 pupils looks really flat. The image looks double, like its jumped partway through the 26x2.5 INT

off source luci1..051, 0.5"

The second on source does not show the same jump, luci1..052-00.53 (0.4" seeing)

luci..54 offsource, image is a bit elongated.

luci..55 & 56 getting 3.2-3.3 pixel FWHM and nice round images with ESM

Script ran with luci on offs. luci1...0059. Last image is an off source so loop is paused. We will try anand send preset from here to see what happens.

2:56 Rerunning script to see about settings after Pause, and verify stability (check for any more jumps). 0.57" on the WFS, cirrus.

luci1..0060 is first image after closing ESM with same gains as first run through. OCAM estimated magnitude 11.9. 3.1 pixels = 0.37"

luci1..0061 offsource

luci1..0062-063 on source.

3:06 Sending preset to M34 for bin2 ESM test. 0.6" seeing on. M34_R12pt4_ESM_L1.xml

OCAM Estimate Mag 13.9, Gain TT:3, HO1:1. Camera lens is railed. This loaded as bin1 since we do not do check flux.

luci1.20211202.0064 on source. Image jump again on first image. Its almost like the image is starting before settling, although Doug is running in intervention mode so that seems unlikely

3:20 sending preset with AOref mag of 12.51 to set it as a bin2 ESM. For somereason it still set it up as bin1??

3:26 send preset with AOref magnitude set to 13.0. The ACE table clearly has 12.5-14.5mag as bin 2 but it keeps setting up as bin1...

8:32 Giving up on AO for now. Reauthorizing as LUCI2 only and David is pointing and collimating.

3:47 Sending StonA2020 rot ZP script. Seeing 0.57" on the Guider. thin cirrus.

XShift 1.84720, -2.42194, Angle -0.05739, so thats as good as one can ask for. Confirmatory image looks good.

4:02 Sending Telluric script for ZP check.
Offset to slit -4.7933, -7.2165 luci2.20211202.0007 acquisition

Bottom left: luci2...0008, top left: luci2...0009, down luci2.202112.0008 (bottom slit is misaligned), top right luci2...011

4:15 Reconfig to the LBC's

-- BarryRothberg - 01 Dec 2021
Topic revision: r5 - 02 Dec 2021, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback